1-Question 28 3.03 pts Convert the radian measure to degrees. Give answer using decimal degrees to the nearest hundredth. Use 3

2-Question 29 3.03 pts Solve the problem. A wheel is rotating at 7 radians/sec, and the wheel has a 44-inch diameter. To the ne
Answer #1

1. Radian=2

1 radian=180 /π degree

2radian=360/3.1416 degree

=114.59⁰

2.

Circumference=πd=44π inches

Wheel is rotating at 7 radian/ sec

Every revolution is of 2π radians

So 7/2π revolution/sec==3.5/π revolution/sec

Speed=44π*3.5/π inches/sec

= 44*3.5*60/12 ft/min

=770ft/min